kayla233:

Tell whether each function is quadratic. Explain. (0,6) (1,12) (2,20) (3,30)

darkknight:

alright so to find whether a function is quadratic or not, u have to do a whole bunch of junk u only need 3 points to determine if a function is quadratic or not \[ax^2+bx+c =y\] so if we take the points (0, 6) (1, 12) and (2,20) \[a(0)^2+b(0)+c = 6\] \[a(1)^2+b(1)+c = 12\] \[a(2^2)+b(2)+c = 20\]

100kpimpcess:

@Laylalyssa

darkknight:

from here you know \[c = 6\] \[a+b+c = 12\] and \[4a+2b+c = 20\] you can substitute the value for c in the other 2 equations as well

darkknight:

\[a+b+6=12\] \[4a+2b+6=20\] _________________________________________ \[a+b=6\] \[4a+2b=14\] We would want to eliminate one variable and solve for the other, think you can do it from here? if u get a = 0 or the equations do not match the function is not quadratic

kayla233:

i think i got it now, thanks

darkknight:

np